OpenStudy (berrymox):

Find the derivative of: sqrt(3t+1) Using h-> 0

OpenStudy (berrymox):

I messed up somehow and got 3. The answer is: 3 ---------- 2 sqrt(x+1)

OpenStudy (phi):

are you supposed to use \[ \lim_{h \rightarrow 0} \frac{ \sqrt{3(t+h)+1} -\sqrt{3t+1} }{h}\]

OpenStudy (berrymox):

Yes. thus far I got: 3 ----------------------- sqrt(3t+1) + sqrt(3x+1)

OpenStudy (phi):

if so, the "trick" is to multiply top and bottom by the conjugate (i.e. same expression but with a plus sign in-between the terms)

OpenStudy (berrymox):

Not sure if I made a mistake at this point.

OpenStudy (phi):

that bottom looks wrong. I would expect it to be \[ \sqrt{3(t+h)+1} +\sqrt{3t+1} \]

OpenStudy (phi):

you seem to be flipping between using "t" and "x" as the variable. but after multiplying top and bottom by the conjugate, and simplifying you have \[ \lim_{h \rightarrow 0} \frac{ 3}{ \sqrt{3(t+h)+1} +\sqrt{3t+1} } \]

OpenStudy (berrymox):

I have 7 min to show my work then go to class

OpenStudy (phi):

the last step is to let h go to zero. that makes the bottom 2 * sqrt(stuff)

OpenStudy (phi):

\[ \lim_{h \rightarrow 0} \frac{ 3}{ \sqrt{3(t+h)+1} +\sqrt{3t+1} } \\ \frac{ 3}{ \sqrt{3t+1} +\sqrt{3t+1} } \]

OpenStudy (berrymox):

thank you